# Continuous Diffusion for Mixed-Type Tabular Data

### v3

@@ Appendix I Machine Learning Efficiency Models
adult $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
bank $0.002_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.040_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.030_{\pm 0.001}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$
beijing $0.002_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.058_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$
churn $0.006_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.027_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.048_{\pm 0.002}$ $0.013_{\pm 0.002}$ $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.006_{\pm 0.001}$
covertype - $0.006_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.015_{\pm 0.000}$ $0.012_{\pm 0.000}$
default $0.002_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
diabetes $0.004_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.042_{\pm 0.000}$
lending $0.006_{\pm 0.000}$ $0.013_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.425_{\pm 0.001}$ $0.050_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.010_{\pm 0.000}$ $0.011_{\pm 0.000}$
news $0.007_{\pm 0.000}$ $0.024_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.078_{\pm 0.001}$ $0.030_{\pm 0.000}$ $0.025_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.008_{\pm 0.000}$
nmes $0.005_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.007_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$
@@
acsincome - $0.808_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.985_{\pm 0.000}$ $0.825_{\pm 0.002}$ $0.680_{\pm 0.002}$ $0.540_{\pm 0.003}$ $0.532_{\pm 0.004}$ $0.526_{\pm 0.002}$
adult $0.320_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.967_{\pm 0.001}$ $0.590_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.630_{\pm 0.003}$ $0.604_{\pm 0.002}$ $0.588_{\pm 0.002}$ $0.591_{\pm 0.005}$
bank $0.633_{\pm 0.008}$ $0.955_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.988_{\pm 0.001}$ $0.783_{\pm 0.003}$ $1.000_{\pm 0.000}$ $0.843_{\pm 0.002}$ $0.795_{\pm 0.003}$ $0.739_{\pm 0.003}$ $0.694_{\pm 0.006}$
beijing $0.976_{\pm 0.001}$ $0.995_{\pm 0.000}$ $0.998_{\pm 0.000}$ $0.993_{\pm 0.001}$ $0.966_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.966_{\pm 0.001}$ $0.951_{\pm 0.002}$ $0.949_{\pm 0.001}$ $0.947_{\pm 0.002}$
churn $0.339_{\pm 0.020}$ $0.853_{\pm 0.002}$ $0.945_{\pm 0.006}$ $0.843_{\pm 0.011}$ $0.561_{\pm 0.005}$ $0.730_{\pm 0.012}$ $0.865_{\pm 0.012}$ $0.621_{\pm 0.016}$ $0.533_{\pm 0.007}$ $0.544_{\pm 0.031}$
covertype - $0.945_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.989_{\pm 0.001}$ $0.586_{\pm 0.002}$ $0.900_{\pm 0.002}$ $0.979_{\pm 0.001}$ $0.991_{\pm 0.001}$ $0.992_{\pm 0.001}$ $0.991_{\pm 0.001}$
default $0.493_{\pm 0.009}$ $0.991_{\pm 0.001}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.821_{\pm 0.002}$ $0.995_{\pm 0.000}$ $0.902_{\pm 0.001}$ $0.827_{\pm 0.004}$ $0.802_{\pm 0.003}$ $0.871_{\pm 0.001}$
diabetes $0.367_{\pm 0.001}$ $0.854_{\pm 0.002}$ $0.935_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.940_{\pm 0.001}$ $0.858_{\pm 0.001}$ $0.780_{\pm 0.002}$ $0.866_{\pm 0.002}$
lending $0.926_{\pm 0.004}$ $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.995_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.998_{\pm 0.001}$ $0.955_{\pm 0.006}$ $0.954_{\pm 0.009}$ $0.961_{\pm 0.004}$
news $0.993_{\pm 0.001}$ $0.998_{\pm 0.000}$ $1.000_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.966_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.999_{\pm 0.000}$ $0.973_{\pm 0.001}$ $0.953_{\pm 0.001}$ $0.977_{\pm 0.001}$
nmes $0.905_{\pm 0.007}$ $0.987_{\pm 0.002}$ $0.992_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.650_{\pm 0.014}$ $0.988_{\pm 0.000}$ $0.841_{\pm 0.008}$ $0.636_{\pm 0.008}$ $0.623_{\pm 0.008}$ $0.642_{\pm 0.010}$
@@
acsincome $7.673_{\pm 0.017}$ $8.637_{\pm 0.027}$ $10.758_{\pm 0.054}$ $6.652_{\pm 0.032}$ $10.877_{\pm 0.092}$ $10.305_{\pm 0.073}$ $8.346_{\pm 0.056}$ $8.322_{\pm 0.047}$ $8.349_{\pm 0.033}$
adult $1.870_{\pm 0.000}$ $1.371_{\pm 0.018}$ $2.523_{\pm 0.012}$ $5.012_{\pm 0.028}$ $2.227_{\pm 0.013}$ $1.647_{\pm 0.009}$ $2.735_{\pm 0.028}$ $2.341_{\pm 0.013}$ $1.112_{\pm 0.019}$ $1.231_{\pm 0.011}$ $1.294_{\pm 0.009}$
bank $2.369_{\pm 0.000}$ $1.369_{\pm 0.011}$ $3.025_{\pm 0.017}$ $3.840_{\pm 0.014}$ $3.136_{\pm 0.007}$ $2.327_{\pm 0.010}$ $3.062_{\pm 0.012}$ $2.973_{\pm 0.012}$ $1.828_{\pm 0.008}$ $1.943_{\pm 0.007}$ $2.062_{\pm 0.008}$
beijing $0.385_{\pm 0.000}$ $0.139_{\pm 0.003}$ $0.735_{\pm 0.003}$ $1.004_{\pm 0.006}$ $0.926_{\pm 0.003}$ $0.739_{\pm 0.006}$ $0.610_{\pm 0.002}$ $0.626_{\pm 0.001}$ $0.490_{\pm 0.002}$ $0.489_{\pm 0.001}$ $0.477_{\pm 0.002}$
churn $0.347_{\pm 0.000}$ $0.232_{\pm 0.028}$ $1.136_{\pm 0.015}$ $1.804_{\pm 0.036}$ $1.146_{\pm 0.039}$ $0.342_{\pm 0.031}$ $0.852_{\pm 0.016}$ $1.130_{\pm 0.018}$ $0.332_{\pm 0.021}$ $0.274_{\pm 0.021}$ $0.276_{\pm 0.012}$
covertype $0.529_{\pm 0.001}$ $1.741_{\pm 0.011}$ $5.773_{\pm 0.017}$ $3.173_{\pm 0.013}$ $0.889_{\pm 0.007}$ $1.508_{\pm 0.020}$ $3.086_{\pm 0.009}$ $2.297_{\pm 0.026}$ $2.209_{\pm 0.022}$ $2.252_{\pm 0.013}$
default $1.812_{\pm 0.000}$ $1.032_{\pm 0.010}$ $3.095_{\pm 0.026}$ $5.880_{\pm 0.020}$ $3.216_{\pm 0.013}$ $1.422_{\pm 0.013}$ $2.593_{\pm 0.020}$ $2.603_{\pm 0.018}$ $1.127_{\pm 0.028}$ $1.269_{\pm 0.014}$ $1.253_{\pm 0.012}$
diabetes $15.608_{\pm 0.055}$ $13.909_{\pm 0.050}$ $17.736_{\pm 0.107}$ $21.935_{\pm 0.046}$ $8.214_{\pm 0.022}$ $28.955_{\pm 0.060}$ $15.279_{\pm 0.026}$ $15.126_{\pm 0.058}$ $15.350_{\pm 0.059}$
lending $11.184_{\pm 0.000}$ $17.752_{\pm 0.143}$ $17.776_{\pm 0.132}$ $20.239_{\pm 0.222}$ $10.688_{\pm 0.025}$ $12.537_{\pm 0.076}$ $16.222_{\pm 0.092}$ $13.775_{\pm 0.147}$ $14.162_{\pm 0.188}$ $13.966_{\pm 0.282}$
news $3.615_{\pm 0.000}$ $3.553_{\pm 0.134}$ $6.147_{\pm 0.010}$ $4.789_{\pm 0.005}$ $5.821_{\pm 0.003}$ $4.960_{\pm 0.006}$ $4.661_{\pm 0.023}$ $5.351_{\pm 0.008}$ $3.635_{\pm 0.004}$ $3.687_{\pm 0.006}$ $3.749_{\pm 0.048}$
nmes $1.931_{\pm 0.000}$ $1.394_{\pm 0.019}$ $2.203_{\pm 0.028}$ $2.971_{\pm 0.008}$ $1.710_{\pm 0.019}$ $0.891_{\pm 0.033}$ $1.231_{\pm 0.024}$ $2.260_{\pm 0.034}$ $0.664_{\pm 0.029}$ $0.710_{\pm 0.032}$ $0.771_{\pm 0.023}$
@@
adult $0.797_{\pm 0.000}$ $0.784_{\pm 0.001}$ $0.769_{\pm 0.002}$ $0.647_{\pm 0.015}$ $0.756_{\pm 0.002}$ $0.787_{\pm 0.001}$ $0.745_{\pm 0.004}$ $0.782_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$
bank $0.745_{\pm 0.002}$ $0.740_{\pm 0.004}$ $0.682_{\pm 0.006}$ $0.680_{\pm 0.006}$ $0.629_{\pm 0.006}$ $0.720_{\pm 0.006}$ $0.673_{\pm 0.006}$ $0.711_{\pm 0.007}$ $0.776_{\pm 0.003}$ $0.767_{\pm 0.004}$ $0.737_{\pm 0.004}$
churn $0.873_{\pm 0.003}$ $0.865_{\pm 0.008}$ $0.780_{\pm 0.015}$ $0.761_{\pm 0.009}$ $0.802_{\pm 0.017}$ $0.857_{\pm 0.007}$ $0.865_{\pm 0.008}$ $0.771_{\pm 0.014}$ $0.854_{\pm 0.011}$ $0.852_{\pm 0.006}$ $0.845_{\pm 0.011}$
covertype $0.817_{\pm 0.001}$ $0.783_{\pm 0.001}$ $0.442_{\pm 0.008}$ $0.711_{\pm 0.002}$ $0.799_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.614_{\pm 0.015}$ $0.734_{\pm 0.002}$ $0.754_{\pm 0.001}$ $0.722_{\pm 0.002}$
default $0.674_{\pm 0.001}$ $0.677_{\pm 0.001}$ $0.627_{\pm 0.003}$ $0.686_{\pm 0.002}$ $0.632_{\pm 0.007}$ $0.678_{\pm 0.002}$ $0.638_{\pm 0.008}$ $0.496_{\pm 0.009}$ $0.670_{\pm 0.002}$ $0.671_{\pm 0.001}$ $0.673_{\pm 0.003}$
diabetes $0.621_{\pm 0.002}$ $0.615_{\pm 0.002}$ $0.572_{\pm 0.005}$ $0.557_{\pm 0.004}$ $0.553_{\pm 0.003}$ $0.560_{\pm 0.006}$ $0.617_{\pm 0.002}$ $0.617_{\pm 0.002}$ $0.611_{\pm 0.002}$
@@
adult $0.915_{\pm 0.000}$ $0.906_{\pm 0.001}$ $0.901_{\pm 0.000}$ $0.836_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.908_{\pm 0.000}$ $0.880_{\pm 0.005}$ $0.906_{\pm 0.001}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$ $0.909_{\pm 0.000}$
bank $0.947_{\pm 0.000}$ $0.943_{\pm 0.001}$ $0.938_{\pm 0.001}$ $0.934_{\pm 0.003}$ $0.830_{\pm 0.020}$ $0.940_{\pm 0.005}$ $0.929_{\pm 0.005}$ $0.939_{\pm 0.003}$ $0.945_{\pm 0.000}$ $0.945_{\pm 0.001}$ $0.943_{\pm 0.004}$
churn $0.964_{\pm 0.001}$ $0.961_{\pm 0.002}$ $0.939_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.948_{\pm 0.004}$ $0.957_{\pm 0.004}$ $0.961_{\pm 0.001}$ $0.919_{\pm 0.006}$ $0.962_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.959_{\pm 0.003}$
covertype $0.892_{\pm 0.000}$ $0.860_{\pm 0.001}$ $0.677_{\pm 0.007}$ $0.777_{\pm 0.001}$ $0.876_{\pm 0.000}$ $0.845_{\pm 0.001}$ $0.671_{\pm 0.013}$ $0.816_{\pm 0.002}$ $0.828_{\pm 0.001}$ $0.802_{\pm 0.002}$
default $0.768_{\pm 0.000}$ $0.759_{\pm 0.003}$ $0.754_{\pm 0.002}$ $0.744_{\pm 0.002}$ $0.751_{\pm 0.004}$ $0.763_{\pm 0.002}$ $0.739_{\pm 0.008}$ $0.746_{\pm 0.011}$ $0.762_{\pm 0.003}$ $0.765_{\pm 0.002}$ $0.765_{\pm 0.002}$
diabetes $0.693_{\pm 0.001}$ $0.679_{\pm 0.001}$ $0.669_{\pm 0.002}$ $0.626_{\pm 0.003}$ $0.592_{\pm 0.002}$ $0.645_{\pm 0.002}$ $0.675_{\pm 0.001}$ $0.673_{\pm 0.001}$ $0.667_{\pm 0.001}$

### v4

@@ Appendix I Tuning of the Detection Model
adult $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$
bank $0.002_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.040_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.030_{\pm 0.001}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.003_{\pm 0.001}$ $0.007_{\pm 0.001}$
beijing $0.002_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.030_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$
churn $0.006_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.027_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.007_{\pm 0.002}$ $0.048_{\pm 0.002}$ $0.013_{\pm 0.002}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$
covertype - $0.006_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.013_{\pm 0.000}$
default $0.002_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$
diabetes $0.004_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.042_{\pm 0.000}$ $0.034_{\pm 0.000}$ $0.041_{\pm 0.000}$
lending $0.006_{\pm 0.000}$ $0.013_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.410_{\pm 0.001}$ $0.053_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.011_{\pm 0.000}$
news $0.007_{\pm 0.000}$ $0.024_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.033_{\pm 0.001}$ $0.030_{\pm 0.000}$ $0.029_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.008_{\pm 0.000}$
nmes $0.005_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.007_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.000}$
@@
acsincome - $0.808_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.985_{\pm 0.000}$ $0.825_{\pm 0.002}$ $0.688_{\pm 0.003}$ $0.529_{\pm 0.005}$ $0.527_{\pm 0.002}$ $0.528_{\pm 0.002}$
adult $0.687_{\pm 0.003}$ $0.889_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.967_{\pm 0.001}$ $0.594_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.641_{\pm 0.003}$ $0.621_{\pm 0.002}$ $0.590_{\pm 0.004}$ $0.605_{\pm 0.004}$
bank $0.839_{\pm 0.003}$ $0.955_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.988_{\pm 0.001}$ $0.781_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.853_{\pm 0.003}$ $0.808_{\pm 0.003}$ $0.701_{\pm 0.005}$ $0.835_{\pm 0.001}$
beijing $0.938_{\pm 0.002}$ $0.989_{\pm 0.002}$ $0.996_{\pm 0.001}$ $0.995_{\pm 0.001}$ $0.738_{\pm 0.004}$ $0.989_{\pm 0.001}$ $0.723_{\pm 0.003}$ $0.574_{\pm 0.003}$ $0.620_{\pm 0.005}$ $0.614_{\pm 0.003}$
churn $0.567_{\pm 0.015}$ $0.853_{\pm 0.002}$ $0.945_{\pm 0.006}$ $0.843_{\pm 0.011}$ $0.556_{\pm 0.013}$ $0.730_{\pm 0.012}$ $0.859_{\pm 0.005}$ $0.614_{\pm 0.016}$ $0.541_{\pm 0.008}$ $0.639_{\pm 0.011}$
covertype - $0.945_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.989_{\pm 0.001}$ $0.584_{\pm 0.002}$ $0.900_{\pm 0.002}$ $0.991_{\pm 0.000}$ $0.989_{\pm 0.000}$ $0.981_{\pm 0.000}$ $0.989_{\pm 0.001}$
default $0.928_{\pm 0.004}$ $0.991_{\pm 0.001}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.827_{\pm 0.005}$ $0.995_{\pm 0.000}$ $0.914_{\pm 0.002}$ $0.823_{\pm 0.001}$ $0.793_{\pm 0.005}$ $0.827_{\pm 0.003}$
diabetes $0.726_{\pm 0.001}$ $0.854_{\pm 0.002}$ $0.935_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.946_{\pm 0.001}$ $0.864_{\pm 0.002}$ $0.837_{\pm 0.001}$ $0.862_{\pm 0.001}$
lending $0.966_{\pm 0.003}$ $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.995_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.998_{\pm 0.000}$ $0.959_{\pm 0.009}$ $0.955_{\pm 0.003}$ $0.960_{\pm 0.005}$
news $0.998_{\pm 0.000}$ $0.998_{\pm 0.000}$ $1.000_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.974_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.999_{\pm 0.000}$ $0.947_{\pm 0.002}$ $0.950_{\pm 0.002}$ $0.972_{\pm 0.002}$
nmes $0.926_{\pm 0.007}$ $0.987_{\pm 0.002}$ $0.992_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.652_{\pm 0.011}$ $0.988_{\pm 0.000}$ $0.829_{\pm 0.010}$ $0.646_{\pm 0.005}$ $0.633_{\pm 0.010}$ $0.636_{\pm 0.010}$
@@
acsincome $7.673_{\pm 0.017}$ $8.637_{\pm 0.027}$ $10.758_{\pm 0.054}$ $6.652_{\pm 0.032}$ $10.877_{\pm 0.092}$ $10.797_{\pm 0.101}$ $8.337_{\pm 0.053}$ $8.397_{\pm 0.062}$ $8.344_{\pm 0.027}$
adult $1.870_{\pm 0.000}$ $1.371_{\pm 0.018}$ $2.523_{\pm 0.012}$ $5.012_{\pm 0.028}$ $2.227_{\pm 0.013}$ $1.656_{\pm 0.008}$ $2.735_{\pm 0.028}$ $2.408_{\pm 0.031}$ $1.138_{\pm 0.015}$ $1.327_{\pm 0.015}$ $1.334_{\pm 0.012}$
bank $2.369_{\pm 0.000}$ $1.369_{\pm 0.011}$ $3.025_{\pm 0.017}$ $3.840_{\pm 0.014}$ $3.136_{\pm 0.007}$ $2.211_{\pm 0.011}$ $3.062_{\pm 0.012}$ $3.022_{\pm 0.009}$ $1.749_{\pm 0.008}$ $1.913_{\pm 0.006}$ $1.997_{\pm 0.008}$
beijing $0.385_{\pm 0.000}$ $0.139_{\pm 0.003}$ $0.731_{\pm 0.003}$ $0.800_{\pm 0.002}$ $0.724_{\pm 0.005}$ $0.639_{\pm 0.002}$ $0.588_{\pm 0.003}$ $0.633_{\pm 0.002}$ $0.474_{\pm 0.002}$ $0.473_{\pm 0.002}$ $0.469_{\pm 0.001}$
churn $0.347_{\pm 0.000}$ $0.232_{\pm 0.028}$ $1.136_{\pm 0.015}$ $1.804_{\pm 0.036}$ $1.146_{\pm 0.039}$ $0.368_{\pm 0.044}$ $0.852_{\pm 0.016}$ $1.209_{\pm 0.010}$ $0.329_{\pm 0.010}$ $0.278_{\pm 0.012}$ $0.350_{\pm 0.023}$
covertype $0.529_{\pm 0.001}$ $1.741_{\pm 0.011}$ $5.773_{\pm 0.017}$ $3.173_{\pm 0.013}$ $0.877_{\pm 0.008}$ $1.508_{\pm 0.020}$ $3.033_{\pm 0.012}$ $1.825_{\pm 0.016}$ $1.594_{\pm 0.009}$ $1.805_{\pm 0.009}$
default $1.812_{\pm 0.000}$ $1.032_{\pm 0.010}$ $3.095_{\pm 0.026}$ $5.880_{\pm 0.020}$ $3.216_{\pm 0.013}$ $1.437_{\pm 0.020}$ $2.593_{\pm 0.020}$ $2.801_{\pm 0.032}$ $1.192_{\pm 0.022}$ $1.355_{\pm 0.017}$ $1.352_{\pm 0.016}$
diabetes $15.608_{\pm 0.055}$ $13.909_{\pm 0.050}$ $17.736_{\pm 0.107}$ $21.935_{\pm 0.046}$ $8.214_{\pm 0.022}$ $28.794_{\pm 0.054}$ $14.356_{\pm 0.050}$ $14.468_{\pm 0.028}$ $14.866_{\pm 0.033}$
lending $11.184_{\pm 0.000}$ $17.752_{\pm 0.143}$ $17.776_{\pm 0.132}$ $20.239_{\pm 0.222}$ $10.688_{\pm 0.025}$ $14.310_{\pm 0.093}$ $16.239_{\pm 0.052}$ $14.958_{\pm 0.292}$ $14.962_{\pm 0.090}$ $14.146_{\pm 0.240}$
news $3.615_{\pm 0.000}$ $3.553_{\pm 0.134}$ $6.147_{\pm 0.010}$ $4.789_{\pm 0.005}$ $5.821_{\pm 0.003}$ $4.358_{\pm 0.013}$ $4.661_{\pm 0.023}$ $5.410_{\pm 0.005}$ $3.615_{\pm 0.009}$ $3.676_{\pm 0.008}$ $3.736_{\pm 0.094}$
nmes $1.931_{\pm 0.000}$ $1.394_{\pm 0.019}$ $2.203_{\pm 0.028}$ $2.971_{\pm 0.008}$ $1.710_{\pm 0.019}$ $0.890_{\pm 0.027}$ $1.231_{\pm 0.024}$ $2.105_{\pm 0.022}$ $0.801_{\pm 0.017}$ $0.780_{\pm 0.031}$ $0.803_{\pm 0.017}$
@@
adult $0.797_{\pm 0.000}$ $0.784_{\pm 0.001}$ $0.769_{\pm 0.002}$ $0.647_{\pm 0.015}$ $0.756_{\pm 0.002}$ $0.788_{\pm 0.001}$ $0.745_{\pm 0.004}$ $0.782_{\pm 0.002}$ $0.788_{\pm 0.001}$ $0.789_{\pm 0.002}$ $0.789_{\pm 0.002}$
bank $0.745_{\pm 0.002}$ $0.740_{\pm 0.004}$ $0.682_{\pm 0.006}$ $0.680_{\pm 0.006}$ $0.629_{\pm 0.006}$ $0.744_{\pm 0.005}$ $0.673_{\pm 0.006}$ $0.661_{\pm 0.008}$ $0.782_{\pm 0.003}$ $0.766_{\pm 0.005}$ $0.747_{\pm 0.004}$
churn $0.873_{\pm 0.003}$ $0.865_{\pm 0.008}$ $0.780_{\pm 0.015}$ $0.761_{\pm 0.009}$ $0.802_{\pm 0.017}$ $0.855_{\pm 0.012}$ $0.865_{\pm 0.008}$ $0.748_{\pm 0.015}$ $0.859_{\pm 0.007}$ $0.863_{\pm 0.007}$ $0.860_{\pm 0.007}$
covertype $0.817_{\pm 0.001}$ $0.783_{\pm 0.001}$ $0.442_{\pm 0.008}$ $0.711_{\pm 0.002}$ $0.799_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.620_{\pm 0.014}$ $0.766_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.765_{\pm 0.001}$
default $0.674_{\pm 0.001}$ $0.677_{\pm 0.001}$ $0.627_{\pm 0.003}$ $0.686_{\pm 0.002}$ $0.632_{\pm 0.007}$ $0.680_{\pm 0.002}$ $0.638_{\pm 0.008}$ $0.485_{\pm 0.016}$ $0.673_{\pm 0.003}$ $0.675_{\pm 0.002}$ $0.677_{\pm 0.002}$
diabetes $0.621_{\pm 0.002}$ $0.615_{\pm 0.002}$ $0.572_{\pm 0.005}$ $0.557_{\pm 0.004}$ $0.553_{\pm 0.003}$ $0.566_{\pm 0.006}$ $0.614_{\pm 0.003}$ $0.619_{\pm 0.002}$ $0.614_{\pm 0.003}$
@@
adult $0.915_{\pm 0.000}$ $0.906_{\pm 0.001}$ $0.901_{\pm 0.000}$ $0.836_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.909_{\pm 0.000}$ $0.880_{\pm 0.005}$ $0.906_{\pm 0.001}$ $0.909_{\pm 0.000}$ $0.909_{\pm 0.001}$ $0.908_{\pm 0.001}$
bank $0.947_{\pm 0.000}$ $0.943_{\pm 0.001}$ $0.938_{\pm 0.001}$ $0.934_{\pm 0.003}$ $0.830_{\pm 0.020}$ $0.942_{\pm 0.004}$ $0.929_{\pm 0.005}$ $0.922_{\pm 0.006}$ $0.945_{\pm 0.000}$ $0.946_{\pm 0.000}$ $0.944_{\pm 0.003}$
churn $0.964_{\pm 0.001}$ $0.961_{\pm 0.002}$ $0.939_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.948_{\pm 0.004}$ $0.957_{\pm 0.006}$ $0.961_{\pm 0.001}$ $0.911_{\pm 0.013}$ $0.960_{\pm 0.002}$ $0.957_{\pm 0.008}$ $0.960_{\pm 0.003}$
covertype $0.892_{\pm 0.000}$ $0.860_{\pm 0.001}$ $0.677_{\pm 0.007}$ $0.777_{\pm 0.001}$ $0.876_{\pm 0.001}$ $0.845_{\pm 0.001}$ $0.675_{\pm 0.013}$ $0.840_{\pm 0.001}$ $0.844_{\pm 0.000}$ $0.840_{\pm 0.001}$
default $0.768_{\pm 0.000}$ $0.759_{\pm 0.003}$ $0.754_{\pm 0.002}$ $0.744_{\pm 0.002}$ $0.751_{\pm 0.004}$ $0.765_{\pm 0.002}$ $0.739_{\pm 0.008}$ $0.732_{\pm 0.021}$ $0.763_{\pm 0.002}$ $0.764_{\pm 0.002}$ $0.765_{\pm 0.002}$
diabetes $0.693_{\pm 0.001}$ $0.679_{\pm 0.001}$ $0.669_{\pm 0.002}$ $0.626_{\pm 0.003}$ $0.592_{\pm 0.002}$ $0.642_{\pm 0.002}$ $0.671_{\pm 0.002}$ $0.673_{\pm 0.002}$ $0.672_{\pm 0.002}$